OpenStudy (anonymous):

Chapter 5: Indices and Logarithms

OpenStudy (anonymous):

\[If~3^{m+1}=5^{n-1}=15^p,show~that~p(m+n)=(m+1)(n-1)\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (mrnood):

you need 2 concepts here - you can do it with these: First \[\log a^{n} = nlog a\] so \[\log 3^{m+1} = (m+1) \log 3\] Second Log (a*b) = log a + log b so log 15 = log (5*3) = log5 + log 3 Armed with that information - have a go at solving for yourself

OpenStudy (alekos):

come on Laino. what are you doing?